OpenStudy (anonymous):

What are the steps in simplifying 10sqrt(20)?

OpenStudy (aravindg):

hmm iits already in a simplified form if you want you may however write \[\sqrt{20}=\sqrt{10 X 2 }\]

OpenStudy (anonymous):

I am a derp. I wrote the wrong thing... I'm sorry...

OpenStudy (thuru7):

\[10\times \sqrt{20} =10\times \sqrt{4\times5}\] \[= 10\times2\sqrt{5}\] \[=20\sqrt{5}\]